OpenStudy (anonymous):

Find the area of a triangle whose base is 2√3 in. and height is 3√3 in.

OpenStudy (jdoe0001):

area of a triangle is \(\bf \cfrac{1}{2} \times base \times height\)

OpenStudy (anonymous):

Ok but I do not understand how to solve for the square root stuff

OpenStudy (jdoe0001):

you multiply outside with outsides, and insides with insides, for the same roots, in this case both have the same root, root 2

OpenStudy (anonymous):

so 2x3 and 3x3 divided by 2

OpenStudy (jdoe0001):

ahemm, yes, but recall the 3's are inside the root, so you keep the insides, inside the root

OpenStudy (jdoe0001):

\(\bf a\sqrt{x} \times b\sqrt{y} \implies ab\sqrt{xy}\)

OpenStudy (anonymous):

I'm lost

OpenStudy (jdoe0001):

well, just follow as => \(a\sqrt{x} \times b\sqrt{y} \implies ab\sqrt{xy}\)

OpenStudy (jdoe0001):

what would that give you?

OpenStudy (anonymous):

Forget it this is pointless.

OpenStudy (anonymous):

jeez what is \(\sqrt{3}\times \sqrt{3}\) ??

OpenStudy (anonymous):

\(\sqrt3\) is the number whose square is 3, so what do you get when you square it? kind of like "who is buried in grant's tomb?" since \(\sqrt3\) is the number you square to get 3, when you square it, you get \(3\)

OpenStudy (anonymous):

in other words \[2\sqrt3\times 3\sqrt3=2\times 3\times \sqrt3\times \sqrt3=6\times 3\]

OpenStudy (anonymous):

18